OpenStudy (peach3113):

-6n-2n=16 What does n equal?

TheSmartOne (thesmartone):

Hint: -6n - 2n = n(-6-2) = ??

OpenStudy (peach3113):

-8?

OpenStudy (xmissalycatx):

What you can do first is combine like terms on the left side which is -8n, then divide both sides by -8, that way 'n' is by itself. :)

OpenStudy (peach3113):

No it's not. I'm confused. What am I doing wrong?

TheSmartOne (thesmartone):

correct

OpenStudy (peach3113):

But wait.

TheSmartOne (thesmartone):

-8n = 18 now you divide both sides by -8

OpenStudy (peach3113):

Uh.

OpenStudy (peach3113):

Okie-dokie.

TheSmartOne (thesmartone):

\(\sf\large \frac{-8n}{-8} =\frac{16}{-8}\) Simplify that :)

OpenStudy (peach3113):

n=-2

TheSmartOne (thesmartone):

and that's your answer! :D

OpenStudy (peach3113):

Ohhh, okay.

OpenStudy (peach3113):

Thank you.

OpenStudy (peach3113):

@TheSmartOne *hugs you* Now I understand. :D

TheSmartOne (thesmartone):

Anytime! :D